Newcastle Fisheries is planning on restarting the salmon hatchery it began in 1887. The new design will cost $9,500. Annual operating costs will be $550 and the equipment will have a salvage value of $1,000 after 30 years. Annual benefits are estimated to be $2,200; annual dis-benefits, $350. If the MARR is 10%, what is the (closest) IRR?

Sam buys a bond with a face value of $1000 for $900. The bond matures in 10 years and pays a coupon interest at 6% every quarter. What is the annual effective rate?

The cost of servers is $360000 and the life of those servers is 10 years. The servers will reduce annual cost by $75000. What is the payback period?
Answer #1

MARR is the minimum rate of return that is expected to be generated from the investment whereas the IRR is the rate at which the net present value of a project is zero.

Cash-flow at period 0 = -9500
Annual cash-flow = 2200-350-550 = 1300
Salvage value at the end of year 30 = 1000

Year CF
0 -9500
1 1300
2 1300
3 1300
4 1300
5 1300
6 1300
7 1300
8 1300
9 1300
10 1300
11 1300
12 1300
13 1300
14 1300
15 1300
16 1300
17 1300
18 1300
19 1300
20 1300
21 1300
22 1300
23 1300
24 1300
25 1300
26 1300
27 1300
28 1300
29 1300
30 2300
IRR 13.40%

IRR formula = IRR(values)

Annual effective rate = (1+0.06/4)^4 - 1 = 1.0614 - 1 = 6.14%
Since the annual rate of 6% is compounded every quarter, the effective rate will be higher.

Payback period = Total costs/ Annual benefits = 360000/ 75000 = 4.8